OpenStudy (anonymous):

graph V(t)=Cos(t^2-2) in the interval of 0 to pi -find average velocity -find the point of maximum velocity and point of minimum velocity (Plz help i dont know how to do this problem and would appreciate an explanation so i know how to do problems in the future similar to this) (im in physics and nt in calculus yet)

OpenStudy (anonymous):

is this really \(V(t)=\cos(t^2-2)\)??

OpenStudy (anonymous):

max is of course 1, because that is the maximum of cosine, and \(\cos(0)=1\) so you would set \(t^2-2=0\) and find that \(t=\sqrt{2}\)

OpenStudy (anonymous):

minimum value of cosine is -1 and \(\cos(\pi)=-1\) so you would have to solve \(t^2-2=\pi\)

OpenStudy (anonymous):

i guess you get \(t=\sqrt{\pi+2}\)

OpenStudy (anonymous):

graph looks like this http://www.wolframalpha.com/input/?i=cos%28x^2-2%29

OpenStudy (anonymous):

are the minumum/max values the same as the minumum/max velocities?